# Optimizing Product Deduplication in E-Commerce with Multimodal Embeddings

### v1

@@ 1 Introduction
Existing solutions often leverage pretrained models due to their strong capabilities in generating meaningful representations. However, these models produce high-dimensional vectors. For example, our initial implementation using the Google Universal Sentence Encoder Multilingual [1] model generated 512-dimensional embeddings. While effective, it faced scalability issues when we hit indexing 8 million items, causing inference times to rise and operational costs to increase. Such limitations highlight the need for scalable, efficient models tailored to the specific requirements of e-commerce.
To address these bottlenecks in the vector search capabilities, we developed a domain-specific text model, fine-tuned for the linguistic and product-specific nuances of Turkish e-commerce, which generates compact product representation. The model built on BERT [2], incorporating dimensionality reduction and enhanced feature extraction through intermediate transformer layers and convolutional layers. This approach enables us to compress representations to 128 dimensions without significant information loss.
@@
2 Related Works
This section presents a review of related works on feature extraction from text and image data, with a specific focus on applications in the e-commerce domain. We also inspect established approaches to product unification.
2.1 Text Based Feature Extraction
Transformer-based architectures have revolutionized semantic representation techniques through sophisticated contextual embedding models. For the Turkish language, there are notable publicly available pre-trained models which have been specifically trained on Turkish text with transformer-based architectures such as BERT, DistilBERT, and ConvBERT [3, 2, 4, 5], each offering unique advantages in terms of performance and efficiency.
BERT (Bidirectional Encoder Representations from Transformers) captures deep bidirectional context by considering both left and right sentence contexts simultaneously [2]. DistilBERT compresses BERT through knowledge distillation, reducing complexity while retaining comparable performance [4]. ConvBERT enhances local feature extraction and efficiency by integrating span-based dynamic convolutions [5].
Domain-specific language model adaptation has emerged as a critical research area in NLP. Gururangan et al. [6] demonstrated the effectiveness of continuing pretraining on domain-specific data before fine-tuning on labeled tasks. Ethayarajh [7] revealed that higher transformer layers capture increasingly context-dependent semantic features.
2.2 Image Based Feature Extraction
Generating representative feature vectors is an important step in image understanding. Convolutional Neural Networks (CNNs) are a cornerstone field in computer vision and paved the way of extracting image features without relying on hand crafted features [8]. Learned image features can be used in various applications in e-commerce such as image search [9, 10], product recommendation [11].
In order to extract image vectors by capturing maximum possible contextual image representation without an extensive labeling process, unsupervised learning based feature extraction systems are developed [12]. Extracted features on these models result in a large vector which is not suitable for deployment in the real world.
2.3 Foundational Multimodal Models
The rapid advancement of vision-language pre-training (VLP) has led to the development of powerful foundational models that learn joint representations of text and images. Contrastive Language-Image Pre-training (CLIP) is a prominent example, employing a dual-encoder architecture to align a vision transformer and a text transformer in a shared latent space via a contrastive loss. This approach, trained on a massive 400 million image-text dataset, enables zero-shot classification and cross-modal retrieval by learning a broad, holistic understanding of visual and textual concepts [13]. While highly general, CLIP’s primary focus on image-level understanding can be a limitation for fine-grained tasks, a challenge that has been addressed by subsequent works such as Alpha-CLIP [14], which introduces an auxiliary channel for region-specific attention, and Long-CLIP [15], which extends the text encoder’s token limit.
Building upon this, BLIP-2 introduced an efficient two-stage bootstrapping strategy that leverages frozen, pre-trained image encoders (e.g., from CLIP) and frozen Large Language Models (LLMs) [16]. This method uses a lightweight Querying Transformer (Q-Former) to bridge the modality gap, achieving state-of-the-art performance on tasks like visual question answering with significantly fewer trainable parameters than previous methods. While these models excel at general-purpose tasks, their architectural and training philosophies differ from our domain-specific approach, which is meticulously designed for the operational realities of e-commerce.
2.4 Product Deduplication Approaches
@@
Matching similar products is an important part of a catalog management system in e-commerce. Duplicate products bloat the product catalog and may cause issues on both operational side and user experience. On the operations side, it increases costs due to increased or redundant manual labor for each product listing. On the other hand, users face redundant search results and overall poor user experience. In order to alleviate these issues,various systems [17, 18, 19, 20, 21] are proposed.
3 Theoretical Analysis
@@ 4 Methodology
4.1 Feature Extraction and Dimensionality Reduction
A central challenge in deploying deep models at scale is managing the high dimensionality of representation vectors. Pretrained models like BERTurk [3] and EfficientNetV2 [22] yield 768- and 1792-dimensional vectors, respectively. These vectors are expensive to store, index, and search in large catalogs.
@@ 4.1 Feature Extraction and Dimensionality Reduction
4.1.1 Text Feature Models
We evaluated several open-source models to generate initial text representations. Among Turkish pretrained language models, as shown in Table I, BERTurk achieved the best performance on our masked language modeling (MLM) task, outperforming ConvBERTurk and DistilBERTurk [3].
@@ 4.1.1 Text Feature Models
To enhance information density and capture lower-level text features, intermediate BERT layer outputs are combined and passed to the compression step (see Figure 1). Non-consecutive intermediate layers were selected, stacked, and processed through convolutional layers before being incorporated similarly to the original BERT output. This modification enabled the model to leverage early Transformer representations, improving its ability to extract richer e-commerce text features.
4.1.2 Visual Feature Models
For the image representations, DeiT [23], Swin Transformer [24], and LeViT [25] were benchmarked on the masked autoencoding (MAE) task. DeiT achieved the highest Multi-scale Structural Similarity Index Measure (MS-SSIM) [26], followed by Swin and LeViT. Results are shared in Table II.
@@ 4.1.2 Visual Feature Models
Exploring contextual image understanding, DeiT-based regular AutoEncoders were tested, but concerns over similar pitfalls led to adopting a Masked AutoEncoder (MAE) approach. Unlike traditional AutoEncoders, MAEs reconstruct masked image regions from unmasked patches, enforcing contextual learning (see Figure 2-a).
@@
To address this, we introduced a structured patch selection method (see Figure 2-b):
@@ 4.1.3 Domain Adaptation
4.2 Product Matching Pipeline
Storing, continuously adding new, and updating product vectors necessitate the use of a dedicated vector database (VecDB) that ensures high availability. Conducting searches using a brute-force approach is resource-intensive; thus, most contemporary VecDB solutions incorporate faster approximate nearest neighbor (ANN) search capabilities through various indexing algorithms [27].
For the VecDB setup, Milvus [28] was chosen for its active open-source community support, ease of setup, distributed capabilities, and its ability to execute various indexing and searching algorithms. The selection of indexing algorithms is crucial for optimizing the efficiency and speed of similarity searches. The two primary candidates considered were the clustering-based Inverted File FLAT (IVF_FLAT) [29] and the graph-based Hierarchical Navigable Small World (HNSW) [30]. Quantized indexing algorithms were deliberately excluded to prevent unnecessary loss in recall rate. Initial tests with varying vector counts and dimensions across both algorithms indicated that, given our memory constraints, IVF_FLAT emerged as the more suitable approach, provided that the appropriate cluster count and query units parameters are selected.
@@ 4.2 Product Matching Pipeline
By choosing a conventional classification approach, the decider model achieves efficient inference, better aligns with available data, and reduces operational complexity. This ensures high performance and robustness across diverse product categories, making the model well-suited for real-world applications in product listing management. The final system design can be seen in Figure 3.
Figure 3: Final system design for product deduplication.
5 Experiments & Results
Training of the architecture utilized a curated e-commerce dataset derived from internal product listings, consisting of 13 million products across 4,500 different categories, each represented by its title and the first product image. With a brief study on the data, for the text representations, the number of max tokens was determined to be 32, which covered %95 of the collected data. In addition, it was apparent that for an e-commerce listing title, most important words tend to exist at the beginning of sentences. Which made taking the first 32 tokens of a product title, a feasible approach.
From the 32 tokens fed as input, BERT created $32\times 128$ length representations with the help of a linear layer for dimensionality reduction. Afterwards for each of the masked input tokens, a classifier layer learned to predict the actual token. After the conclusion of the training, CLS token for each sentence was used as the final text representation. The base model trained on the masked language modeling (MLM) task achieved a validation loss of 1.51.
After further examination of the results and requirements regarding the product titles, we recognized that while BERT effectively captures contextual information within sentences, supplementing its representation vector with lower-level information could lead to additional improvements. To achieve this, intermediate transformer layer outputs from BERT—specifically from layers 1, 3, 5, 7, 9, and 11—were extracted alongside the outputs from the last transformer layer. These layers were then stacked and passed through a convolutional chain followed by normalization and activation layers. Incorporating these intermediate layers allowed the convolutional layers to selectively utilize non-contextual or lower-level features from earlier layers if and when they prove to be more beneficial for the task. This enhanced model achieved a lower MLM validation loss of 1.43, indicating a clear improvement in capturing relevant token-level information. In the final part of the architecture, the [CLS] token from the convolutional chain output was compressed into a 128-length representation and subsequently used in downstream tasks.
As a feasibility study, extraction of image feature vectors were initially made with a simple AutoEncoder, utilizing an EfficientNetV2-Large model as the Encoder and series of convolutional layers as a simple decoder. Utilizing a previous experience with such models, recreation losses were calculated using MS-SSIM since it provided a robustness towards various issues that might occur with MSE (Blurring of target, image scale issues). Since the goal of the development is creating a production ready, searchable and indexable representation vector, outputs of the EfficientNet were compressed to 128, which was decided during trials with randomly generated vectors of different dimensions in order to adhere to the production requirements. This vector size made it possible to scalably index and search available product data.
After conclusion of the feasibility study, research on image representation vectors was made. In order to retain as much of the contextual information as possible Masked AutoEncoders (MAE) surfaced as a suitable approach. Contextual information is found necessary because simple reconstruction mostly focuses on colors and shapes while what the object is also important in the E-commerce domain.
Fundamentally, a MAE’s primary objective is to extract information from an image by examining its individual components. This forces the model to “understand” the objects in the image in order to create a complete reconstruction. As a brief summary to inner workings of MAE therefore can be given as follows;
  • Divide image into N patches,
  • Randomly select %X of N, mask selected,
  • Using an image transformer (DeiT), create features from each unmasked patch,
  • Reconstruct masked patches with the information gained from unmasked patches.
By default, most of the successful MAE algorithms [12], create relatively large representation vectors. While this approach proved successful in recreating images, our requirement for compressed representations made this logic impractical. As a solution, an architecture to create smaller representations was constructed. Specifically, the architecture was designed to work with the DeiT backbone, compressing and concatenating the created per patch representations to a singular 128 length vector.
The resulting representation was used as the output from the “forward” path of the model with a single 128 sized representation vector for each image in the batch. This representation is then fed through convolutional layers to be converted into expected dimensions for MAE decoder. Following this, output was further processed within the decoder, together with the masked token ids, unmasked token ids and unmasked patches (same as in normal MAE) to reconstruct the input image.
While providing a robust starting point for contextual image understanding, MAE has an inherent issue in image feature extraction. Since it randomly selects patches, some small differences within images has a possibility to not be represented in the resulting vector. These sort of differences in objects is a common occurrence in the e-commerce domain, hence retention of these differences becomes relatively important when deciding whether two objects match each other or not. Our internal analysis showed that most of the object details exist on the center of the image. In order to leverage this a more strict patch selection logic was set up. Instead of random patch selection, dataloader was modified to divide the image into 9 equal patches and then create 4 patches as follows;
  • Select center patch, for details that should be preserved,
  • 2 random patches to retain the contextual understanding of MAE,
  • Entire resized image, for details that could help information retention.
From this input a full image was then reconstructed. Another variant that selects 3 random patches in addition to the center patch was also trained. However, this variant did not achieve the expected improvement. This resulted in the hypothesis that the overall outline of the object acquired from the resized full image, assists context understanding and image reconstruction.
As opposed to the initial implementation of MAE, since the input patches include the entire image hence did not have the strict “patch ids” required, recreating just the “masked” parts of the image were not immediately feasible. Therefore this iteration of the architecture learned to reconstruct the full image from the compressed vector.
In the matching trials, an issue kept surfacing regarding the object scales in images. Problems occurred when objects did not take the same space within the image, creating empty spaces between objects and the image borders. This caused vectors to be distanced from each other as if they were variants instead of the same product. An example of this phenomenon can be seen in Figure  4.
Initially as a solution, a separate module was added prior to the dataloader as a data preprocessing method to remove “boring” pixels, closing the gap between object and image border. However this approach quickly proved costly for the production environment, given the sheer number of images processed. Instead, in order to increase the robustness of the model, the same boring pixel removal module has been used in the dataloader, tied to a random number generator. Algorithm for which is as follows:
  • If random number is less than the threshold proceed, if not return original image,
  • Convert image to black and white,
  • Find first non-zero pixel coordinates as (x1,y1), (x2,y2), (x3,y3) and (x4,y4),
  • Extract the center box with found coordinates as points on the edge of a bounding box,
  • Create a background image with the colour as the ”boring pixels”,
  • Scale the bounding box randomly,
  • Embed bounding box to the generated background and return as input image.
Figure 5: Input and Recreation Results of the final iteration of Image Model
During training, input image is cleaned from boring pixels and marked as the reconstruction target. If the generated random number is below the threshold, input of the model will be the image generated with the aforementioned algorithm. If not, input image will be the unprocessed original image. Image reconstruction outputs are provided in Figure 5.
@@ 5 Experiments & Results
In the training procedure, AdamW optimizer was utilized with a starting learning rate of 0.0001 then, a cosine scheduler scaled this value down to near ”0”. Image and text models are trained until convergence. While trials with different optimizers and learning rate values are made, this setup proved most efficient, resulting in lowest loss values.
In the culmination of the study, we address the challenge of product matching in the e-commerce domain by leveraging previously explained multimodal vector data. As detailed in the preceding sections, we employed sophisticated vector extraction techniques to generate dense embeddings from both textual and visual inputs. These vectors serve as robust representations capturing the semantic and visual nuances of each product. Recognizing that product similarity can be ambiguous when relying solely on either text or image data, we hypothesized that combining both modalities would provide a more comprehensive basis for accurate matching. While textual embeddings encapsulate descriptive and categorical information, visual embeddings capture design, color, and style attributes. By integrating these two sources of information, our approach aims to mitigate the limitations inherent in single-modal matching and enhance the overall precision of product pair assessments.
To efficiently handle the high-dimensional vector data and facilitate rapid similarity searches, we utilized Milvus, an open-source vector database optimized for scalable similarity search. Milvus enabled us to perform vector-based retrieval operations with low latency, ensuring that our system remains responsive even as the product catalog expands. By indexing the combined text and image vectors in Milvus, we streamlined the retrieval process, allowing us to quickly identify candidate product pairs for further evaluation by the decider model. In order to evaluate the memory usage of Milvus, we indexed various sizes of randomized vectors into a single-server Milvus setup and measured the system’s memory consumption. This approach allowed us to assess how different vector sizes impacted memory usage. The results, presented in Table  III, indicated that indexing tens of millions of vectors would be practical only with a reduced vector size of 128.
TABLE III: RAM Usage of 10 Million Vectors on Milvus

### v2

@@ I Introduction
Existing solutions often leverage pretrained models due to their strong capabilities in generating meaningful representations. However, these models produce high-dimensional vectors. For example, our initial implementation using the Google Universal Sentence Encoder Multilingual [cer2018universal] model generated 512-dimensional embeddings. While effective, it faced scalability issues when we hit indexing 8 million items, causing inference times to rise and operational costs to increase. Such limitations highlight the need for scalable, efficient models tailored to the specific requirements of e-commerce.
To address these bottlenecks in the vector search capabilities, we developed a domain-specific text model, fine-tuned for the linguistic and product-specific nuances of Turkish e-commerce, which generates compact product representation. The model built on BERT [kenton2019bert], incorporating dimensionality reduction and enhanced feature extraction through intermediate transformer layers and convolutional layers. This approach enables us to compress representations to 128 dimensions without significant information loss.
@@
II Related Works
Our work builds upon established research in feature extraction and multimodal learning, which we adapt for the specific operational constraints of a large-scale e-commerce environment.
II-A Text and Image Representation
Transformer-based architectures, particularly BERT and its variants, have become the standard for semantic text representation by capturing deep bidirectional context [kenton2019bert, sanh2019distilbert, jiang2020convbert]. We leverage BERTurk, a model pre-trained for the Turkish language, as our foundation [stefan_schweter_2020_3770924]. Furthermore, the practice of adapting language models to specific domains is critical for capturing the nuances of e-commerce product titles, which often contain domain-specific jargon and structure.
For image understanding, the field has progressed from Convolutional Neural Networks (CNNs) to transformer-based approaches [shaheen2016impact]. Unsupervised methods like Masked Autoencoders (MAEs) are particularly relevant, as they compel a model to learn contextual visual features rather than just surface-level patterns [he2022masked]. This deeper understanding is crucial for accurately differentiating visually similar but distinct products.
II-B Foundational Multimodal Models and Our Approach
The advent of large-scale Vision-Language Models (VLMs) like CLIP and BLIP-2 has enabled powerful joint representations of text and images. Trained on massive web-scale datasets, these models excel at general-purpose, zero-shot tasks by learning a broad, holistic understanding of visual and textual concepts [radford2021learning, li2023blip].
However, this generality comes at a significant operational cost. Their large parameter counts and high-dimensional vector outputs (e.g., 512 dimensions or higher) introduce prohibitive inference latency and memory requirements for real-time industrial applications like large-scale product deduplication. Our work diverges from this philosophy by prioritizing a domain-specific, efficiency-first approach. Instead of leveraging a single, large VLM, we design separate, lightweight models that generate compact, fast, and highly tailored embeddings to solve a targeted operational problem.
II-C Product Deduplication Approaches
@@
Matching similar products is an important part of a catalog management system in e-commerce. Duplicate products bloat the product catalog and may cause issues on both operational side and user experience. On the operations side, it increases costs due to increased or redundant manual labor for each product listing. On the other hand, users face redundant search results and overall poor user experience. In order to alleviate these issues,various systems [li2020deep, tracz-etal-2020-bert, peeters2022supervised, peeters2024entitymatchingusinglarge, herrero2024learning] are proposed.
III Theoretical Analysis
@@ IV Methodology
IV-A Feature Extraction and Dimensionality Reduction
A central challenge in deploying deep models at scale is managing the high dimensionality of representation vectors. Pretrained models like BERTurk [stefan_schweter_2020_3770924] and EfficientNetV2 [DBLP:journals/corr/abs-2104-00298] yield 768- and 1792-dimensional vectors, respectively. These vectors are expensive to store, index, and search in large catalogs.
@@ IV-A Feature Extraction and Dimensionality Reduction
IV-A1 Text Feature Models
We evaluated several open-source models to generate initial text representations. Among Turkish pretrained language models, as shown in Table I, BERTurk achieved the best performance on our masked language modeling (MLM) task, outperforming ConvBERTurk and DistilBERTurk [stefan_schweter_2020_3770924].
@@ IV-A1 Text Feature Models
To enhance information density and capture lower-level text features, intermediate BERT layer outputs are combined and passed to the compression step (see Fig. 1). Non-consecutive intermediate layers were selected, stacked, and processed through convolutional layers before being incorporated similarly to the original BERT output. This modification enabled the model to leverage early Transformer representations, improving its ability to extract richer e-commerce text features.
@@
IV-A2 Visual Feature Models
For the image representations, DeiT [DBLP:journals/corr/abs-2012-12877], Swin Transformer [DBLP:journals/corr/abs-2103-14030], and LeViT [DBLP:journals/corr/abs-2104-01136] were benchmarked on the masked autoencoding (MAE) task. DeiT achieved the highest Multi-scale Structural Similarity Index Measure (MS-SSIM) [wang2003multiscale], followed by Swin and LeViT. Results are shared in Table II.
@@ IV-A2 Visual Feature Models
Exploring contextual image understanding, DeiT-based regular AutoEncoders were tested, but concerns over similar pitfalls led to adopting a Masked AutoEncoder (MAE) approach. Unlike traditional AutoEncoders, MAEs reconstruct masked image regions from unmasked patches, enforcing contextual learning (see Fig. 2-a).
@@
To address this, we introduced a structured patch selection method (see Fig. 2-b):
@@ IV-A3 Domain Adaptation
IV-B Product Matching Pipeline
Storing, continuously adding new, and updating product vectors necessitate the use of a dedicated vector database (VecDB) that ensures high availability. Conducting searches using a brute-force approach is resource-intensive; thus, most contemporary VecDB solutions incorporate faster approximate nearest neighbor (ANN) search capabilities through various indexing algorithms [pan2024survey].
For the VecDB setup, Milvus [10.1145/3448016.3457550] was chosen for its active open-source community support, ease of setup, distributed capabilities, and its ability to execute various indexing and searching algorithms. The selection of indexing algorithms is crucial for optimizing the efficiency and speed of similarity searches. The two primary candidates considered were the clustering-based Inverted File FLAT (IVF_FLAT) [5432202] and the graph-based Hierarchical Navigable Small World (HNSW) [8594636]. Quantized indexing algorithms were deliberately excluded to prevent unnecessary loss in recall rate. Initial tests with varying vector counts and dimensions across both algorithms indicated that, given our memory constraints, IVF_FLAT emerged as the more suitable approach, provided that the appropriate cluster count and query units parameters are selected.
Figure 3: Final system design for product deduplication.
@@ IV-B Product Matching Pipeline
By choosing a conventional classification approach, the decider model achieves efficient inference, better aligns with available data, and reduces operational complexity. This ensures high performance and robustness across diverse product categories, making the model well-suited for real-world applications in product listing management. The final system design can be seen in Fig. 3.
V Experiments & Results
To validate the effectiveness and efficiency of our proposed multimodal deduplication system, we conducted a series of experiments on a large-scale, real-world e-commerce dataset. We evaluated our model’s performance against a strong commercial baseline and analyzed its operational efficiency in terms of inference speed and memory consumption. Furthermore, we present ablation studies to justify our key architectural decisions, including our compact embedding strategy and text feature aggregation method. This section details the experimental setup, presents the main performance results, and concludes with an error analysis to identify areas for future improvement.
V-A Experimental Setup
This section details the experimental setup used to evaluate our system, including the dataset composition, the baseline model for comparison, and the metrics for performance assessment.
V-A1 Dataset
All models were trained and evaluated on a curated internal dataset from the Hepsiburada e-commerce catalog, consisting of 13 million products across 4,500 different categories. The feature extraction models were pre-trained on this large corpus of product titles and their corresponding primary images. The decider model was trained on a set of 350,000 manually-labeled product pairs and tested on a hold-out set of 8,000 pairs, which were labeled by internal catalog experts. While this internal dataset ensures high domain relevance, we acknowledge that evaluation on public e-commerce datasets would further strengthen the generalizability of our findings.
V-A2 Baseline
We compare our system against a black-box, commercially available API-based deduplication service, referred to as ”3rd Party.” This represents a strong, production-level alternative.
V-A3 Metrics
We report Precision, Recall, and F1-score for both ”Match” and ”Not Match” classes. The primary evaluation metric is the macro-average F1-score, which provides a balanced measure of performance across both classes.
V-A4 Training Procedure
@@
For both text and image models, we utilized the AdamW optimizer with a starting learning rate of 0.0001, complemented by a cosine scheduler that scales the rate down toward zero. This setup proved most efficient, resulting in the lowest loss values during training until convergence.
V-B Main Results
The primary evaluation of our system’s performance is presented in Table III. Our model achieves a macro-average F1 score of 0.90, demonstrating a significant and robust improvement over the third-party solution’s score of 0.83. A deeper analysis of the per-class metrics reveals the practical advantages of our approach for the e-commerce domain.
TABLE III: Performance Comparison Results
Model Class Precision Recall F1
Internal Match 0.91 0.85 0.88
3rd Party Match 0.90 0.73 0.81
Internal Not Match 0.88 0.96 0.92
3rd Party Not Match 0.76 0.94 0.84
The most critical improvement is in the ”Match” class, where our system achieves a Recall of 0.85, compared to 0.73 for the baseline. This 12-point increase is highly significant, as it indicates that our system is substantially more effective at its core task: identifying true duplicate products. This directly translates to a cleaner product catalog, reduced customer confusion, and more effective Buy Box management. Although our precision is marginally higher (0.91 vs. 0.90), the substantial gain in recall underscores our model’s superior ability to find duplicates that the other system misses. For the ”Not Match” class, our system maintains excellent performance with a Recall of 0.96 and Precision of 0.88. This demonstrates that our model is highly reliable in distinguishing between different products, minimizing false positives that would otherwise require costly manual review and erode trust in the automated system.
It is important to note the methodology used to generate the test data. The 8,000 labeled pairs were curated from candidates flagged by our system and the third-party solution. Although this introduces a known bias—focusing on ambiguous cases where both systems are active—it serves as a necessary and pragmatic starting point for a direct, head-to-head comparison. It effectively evaluates how each system performs on a challenging set of real-world examples, allowing us to identify clear areas for improvement and demonstrate our system’s superior performance on this competitive ground.
V-C System Efficiency and Architectural Analysis
To validate the system’s practicality for industrial applications and to justify our key design choices, we analyzed its efficiency and the architectural decisions that led to the final model.
V-C1 System Efficiency and Scalability
A core design principle of our system was operational efficiency, ensuring that it could scale cost-effectively to hundreds of millions of products. Our experimental results validate this approach. As detailed in Table IV, the memory footprint for indexing 10 million of our 128-dimensional vectors is a manageable 5.5 GB. This represents a four-fold reduction compared to the 20+ GB that would be needed for more standard 512-dimensional embeddings, a critical advantage for large-scale deployment.
TABLE IV: RAM Usage of 10 Million Vectors on Milvus